Assume a bank offers an effective annual rate of 5.64%. If compounding is quarterly what is the APR? SHOW WORK
Answer #1

Effective annual rate = [(1+ APR/n]^n - 1

here,

Effective annual rate = 5.64%

=>0.0564

APR = to be found out

n = 4....(quarterly)...

now,

0.0564 = [(1+APR/4)^4 - 1

=>(1.0564)^1/4 = (1+APR/4)

=>1.01381123=1+ APR/4

=>0.01381123=APR/4

=>APR=0.05524492

=>APR=5.52%.